Hitchhiker’s Guide: B Ark  IN: Spring 1998

Society of Cleanliness Engineers

It’s time to clean up this town!

The unenlightened call you Janitors, but you are Cleanliness Engineers. Without you the world halt, dragged down by its own muck and dust. Bob Agrajag A bit paranoid, but a great man with a mop.

Society of Cleanliness Engineers  1  Not a Game Item